A 1 hp motor rotates 1200 rpm.How much torque can it exert?
1
Expert's answer
2020-06-01T14:27:26-0400

Toque

"(T) = \frac{Power(HP)\times 5252 }{RPM}"


"(T) = \frac{1 \times 5252 }{1200}"


"(T) = 4.38 lb-ft = 5.94 N-m"